OpenStudy (anonymous):

7th grade math help plzzz

OpenStudy (anonymous):

@sevy_17

OpenStudy (anonymous):

Which equation illustrates the division property of equality?

OpenStudy (anonymous):

@DJ3strella

OpenStudy (anonymous):

@ineedhelp200

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@grray7

OpenStudy (anonymous):

could u help me with this problem???

OpenStudy (anonymous):

Solve for h. V = lwh A. h = Vlw B. h = C. h = D. h = V – l – w

OpenStudy (anonymous):

I think it's the second equation for your first q

OpenStudy (anonymous):

I need help with my 2nd q

OpenStudy (anonymous):

The multiple choice options don't make sense

OpenStudy (anonymous):

A.h = Vlw B. h =V/lw C. h = lw/V D. h = V – l – w

OpenStudy (anonymous):

Ok, your answer is B.

OpenStudy (anonymous):

Which literal equations are equivalent to k= F/x? Choose all answers that are correct. A. F= kx B. x= kf C. F= k/x D. X= F/k

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@heavenfun13

OpenStudy (anonymous):

@workalltogether

OpenStudy (anonymous):

@chris215

OpenStudy (anonymous):

a,d,

OpenStudy (anonymous):

for the firs ? is the second option :)